OpenStudy (anonymous):

Find a polynomial with the lowest degree, with real coefficients whose zeros include, 4,-10, 3+4i I've gone through the problem and i dont have an answer that matches the choices im not sure if i did it wrong or if the paper is incorrect so checking here!

OpenStudy (jdoe0001):

hmm so...hmmm say... if the polynomials real zeros were say 1 and 2 how would you find it then?

OpenStudy (anonymous):

What i was trying in the book was that you basically multiply all the 0's together in a (x-c) format in which I did but I got an incorrect answer.

OpenStudy (jdoe0001):

gimme a sec

OpenStudy (jdoe0001):

so your roots are 4,-10, 3+4i there's a complex one complex roots do not travel all by their lonesome they come along with their "conjugate" buddy so that means, there is a 4th root that is \(\Large 3-4i\)

OpenStudy (jdoe0001):

so gimme a few secs

OpenStudy (jdoe0001):

hold the mayo

OpenStudy (jdoe0001):

\(\bf x\to \begin{cases} 4\\-10\\ 3+4i\\ 3-4i \end{cases}\implies (x-4)(x+10)(x-3-4i)(x-3+4i) \\ \quad \\ (x-4)(x+10)({\color{brown}{ [x-3] }} -4i)({\color{brown}{ [x-3] }}+4i) \\ \quad \\ recall\implies \textit{difference of squares} \\ \quad \\ (a-b)(a+b) = a^2-b^2\qquad \qquad a^2-b^2 = (a-b)(a+b)\qquad thus \\ \quad \\ (x-4)(x+10)({\color{brown}{ [x-3] }}^2 -[4i]^2) \\ \quad \\ \textit{keep in mind that }i^2=-1\qquad thus \\ \quad \\ (x-4)(x+10)([x-3]^2 -[4^2i^2]) \\ \quad \\ (x-4)(x+10)([x-3]^2 -[16{\color{brown}{ (-1)}}]) \\ \quad \\ (x-4)(x+10)([x-3]^2+16)\\ \quad \\ (x-4)(x+10)([x^2-6x+9]+16) \\ \quad \\ (x-4)(x+10)(x^2-6x+25)\) and from there is just a straight multiplication

OpenStudy (anonymous):

Oh wow! that is the right answer, apparently some how I was getting x^2-6x-7 and not + 25 and that screwed me over lol and I did the problem twice honestly im not sure where I went wrong but I'll lo0k over your answer

OpenStudy (anonymous):

Realized where I went wrong, I did not know i^2 was -1 I assumed it wuld just cancel out into 1.

OpenStudy (anonymous):

Thank you!